# XPUTimer: Anomaly Diagnostics for Divergent LLM Training in GPU Clusters of Thousand-Plus Scale

### v1

@@ 6 Evaluation
In this section, we present experiments to demonstrate XPUTimer’s effectiveness from various perspectives. Specifically, we first evaluate XPUTimer’s runtime overhead in terms of latency and memory, showcasing its lightweight nature for runtime anomaly diagnosis. Then, we assess the effectiveness of XPUTimer’s novel diagnostic mechanisms: intra-kernel tracing for communication hang error diagnosis and issue distribution metric for kernel-issue stall diagnosis.
@@ 6.1 Runtime Overhead
We evaluate XPUTimer on four parallel backbones: Megatron [9], FSDP [8], DeepSpeed [10], and TorchRec [13]. Among these, Megatron, FSDP, and DeepSpeed are widely used for LLM training, while TorchRec is employed for training large recommendation models within Ant Group. Four models are benchmarked, spanning language, vision, and recommendation tasks: two large language models (Llama 18B and 70B), one large vision model (Llama Vision 40B), and one recommendation model (DLRM 72M).
@@
The latency overhead experiment is conducted on 1,024 H800 GPUs deployed across 128 servers with RoCE connectivity. Figure 9 presents the latency overhead introduced by XPUTimer with various models, backbones, and number of GPUs. As shown, XPUTimer incurs a latency overhead of $0.43\%$ for three LLM training backbones and $1.02\%$ for TorchRec.
@@
The memory overhead experiment is conducted on two setups, which are 16 A100 GPUs on 2 nodes and 1536 H800 GPUs on 192 nodes. We compare XPUTimer with Torch Full, Torch w/o Stack, Torch w/o Layout&Stack, and XPUTimer. Torch w/o Stack refers to using the PyTorch builtin profiler with stack tracing disabled, while Torch w/o Layout&Stack further disables matrix layout tracing.
[…]
@@
From the above results, we can conclude that, regardless of the GPU scale, parallel backbone, parallel strategy, or model type used, XPUTimer consistently maintains an extremely low runtime overhead in terms of both latency and memory. The lightweight selective tracing facilitates XPUTimer’s deployment within our large training cluster, serving as a diagnostic framework for diverse training jobs.
@@
Figure 10: Memory consumption of dumped logs per GPU per step using the PyTorch profiler and XPUTimer while training a Llama-70B model on 16 A100 GPUs.
6.2 Effectiveness of Intra-kernel Tracing
@@
We evaluate the intra-kernel tracing mechanism on 16 A100 GPUs across two servers with RoCE connectivity. Given that most communication kernels are ring-based, this experiment focuses on evaluating ring-allreduce. Specifically, we customize the training script composed solely of communication kernels, with one GPU intentionally suspended to simulate a hang error caused by communication issues.
@@
Figure 11 illustrates the pinpointing latencies for intra- and inter-server communication. The figure presents the latency results for three communication protocols [48] and cross-node configurations. As shown, XPUTimer requires $29.4\text{\textasciitilde}309.2s$ to detect erroneous GPUs across different scenarios. Among the protocols, XPUTimer performs best when the SIMPLE protocol is used for communication. This is because, with the SIMPLE protocol, XPUTimer only needs to scan the first thread of each thread block to check the steps, whereas the other two protocols require scanning the entire thread block.
@@
When comparing intra-server and inter-server results, XPUTimer performs better when the ring-allreduce operation spans multiple servers. This is because intra-server GPUs are connected via NVLink, whereas inter-server GPUs communicate through NICs. Communication kernels launched over NICs involve fewer thread blocks, as NICs have fewer internal links compared to NVLink. As a result, XPUTimer scans fewer thread blocks for error diagnosis in inter-server scenarios.
@@
In summary, the intra-kernel tracing mechanism can detect erroneous GPUs in a maximum of $309.2s$. Notably, as the complexity of intra-kernel tracing is $O(1)$, these results remain consistent regardless of scale.
@@
Figure 12: Issue distribution across all communication kernels for a Llama-20B model trained with Megatron and 256 GPUs, including the overall CDF and the CDFs for each type of communication kernel, respectively.
@@ 6.3 Effectiveness of Issue Latency Distribution
In this experiment, we evaluate the issue latency distribution using Llama-20B running on 256 H800 GPUs across 32 servers connected via RoCE. Figure 12 illustrates the issue latency distribution for all communication kernels in the Unhealthy-GC, Unhealthy-Sync, and Healthy scenarios. In the Unhealthy-GC scenario, GC is implicitly triggered by the Python runtime. In the Unhealthy-Sync scenario, an unintended GPU synchronization call is added within the transformer block, leading to repetitive GPU synchronizations during the model’s forward pass. In the Healthy scenario, GC is efficiently managed by the parallel backbone, and no unnecessary synchronizations are introduced.
As shown in the figure, the issue latency distribution patterns align with our claim in §5.2.2. The issue latency CDF of a healthy LLM training job increases linearly, whereas the issue latency CDFs for Unhealthy-GC and Unhealthy-Sync exhibit a much steeper rise. This is because the issue latencies of different ranks in the healthy scenario are solely influenced by the collective communication operator, resulting in a uniform distribution. In contrast, in the cases of Unhealthy-GC and Unhealthy-Sync, while some ranks are affected, their latencies become very short due to the delayed start of the issue time.
Since both GC and GPU synchronizations span the entire model forward pass, all communication kernels are affected, as illustrated in Figure 12. Furthermore, each training process triggers GC independently, and the GC operation for a single process is more time-consuming than GPU synchronization. Consequently, the issue latency distribution for Unhealthy-GC is worse than that of Unhealthy-Sync.
@@ 7 Deployment & Case Studies
In this section, we further demonstrate XPUTimer’s effectiveness in diagnosing LLM training anomalies through statistical analysis of its cluster-wide deployment and detailed case studies of various anomalies.
@@ 7.1 Cluster-wide Deployment
XPUTimer has been deployed within a training cluster with 6,000 GPUs for over 8 months. During this period, it is responsible for monitoring, detecting, and diagnosing training jobs for various deep learning models, especially large-scale distributed LLM training. Its diagnostic capabilities enable the algorithm teams, infrastructure team, and operations team to seamlessly enhance model training efficiency and improve training cluster utilization.
Table 2: Typical errors detected by XPUTimer.
@@
Table 2 presents a subset of error anomalies detected by XPUTimer. As shown in the table, XPUTimer effectively identifies OS- and hardware-related anomalies, including crashes, hangs, and slowdowns caused by such issues. While these issues are typically conspicuous and could be detected by existing methods based on noticeable training interruptions [9, 14], XPUTimer’s novelty lies in providing richer error information through techniques like intra-kernel tracing. Such runtime information helps ease and accelerate the attribution process of low-level issues for operations teams.
Moreover, the true value of XPUTimer lies in its ability to detect anomalies caused by both algorithm teams and the infrastructure team. Table 3 summarizes slowdowns diagnosed by XPUTimer using its aggregated metrics. In the table, anomalies newly identified by XPUTimer, as compared to prior works [2], are highlighted in bold. The following subsections provide a detailed discussion of typical cases.
Table 3: Slowdowns diagnosed by XPUTimer, with “Details” showing training job specifics and associated MFU decline.
@@
Backbone migration 1856 GPUs, Llama-80B, $33.3\%\downarrow$
@@
7.2 Case-1: Towards Stall-free Kernel Issuing
@@
Kernel-issue stalls are among the most frequent causes of slowdowns encountered in a training cluster not dedicated exclusively to a single pre-training task. While Python runtime GC is well-known and now carefully managed by the parallel backbone in most cases, most encountered kernel-issue stalls arise from code introduced by algorithm teams to enhance LLM performance in downstream tasks.
@@
A typical case encountered by XPUTimer is a training job of Llama20B running on 256 H800 GPUs. The developer from the algorithm team mistakenly enables the timer provided by Megatron for performance profiling of several key code segments. This profiling incurs kernel-issue stalls because it requires GPU synchronizations to obtain accurate timestamps.
Although no significant slowdown was observed in training throughput, XPUTimer successfully detects the abnormal issue latency distribution. After removing these unnecessary synchronizations by disabling the timer, the MFU of the training job improves from $41.4\%$ to $42.5\%$, representing a $2.66\%$ increase. While this slowdown is much smaller than that caused by Python runtime GC (typically exceeding $10\%$), XPUTimer could still uncover it, ensuring the training job’s performance.
@@
In addition to the above two cases, XPUTimer also detects other kernel-issue stalls, such as unnecessary package version checking, frequent CUDA memory management within the PyTorch runtime, and others. These cases are listed in Table 3.
7.3 Case-2: Migration between Backbones
Different parallel backbones are suited to varying hardware conditions. While FSDP typically delivers ease of use and good efficiency on fewer than 1000 GPUs or relatively short sequences for LLMs (e.g., 4k), Megatron demonstrates superior scalability as the scale increases beyond 1000 GPUs or when handling extremely large sequences (e.g., 64k). In this context, algorithm teams may migrate an LLM between backbones to meet their specific demands. However, this migration process can potentially introduce anomalies.
@@ 7.3 Case-2: Migration between Backbones
A typical scenario encountered by XPUTimer is an anomalous MFU decline when migrating a Llama-like 80B model from FSDP (1888 H800 GPUs) to Megatron (1586 H800 GPUs with a data-parallel degree of 58, pipeline-parallel degree of 8, and tensor-parallel degree of 4). This anomaly specifically stems from a matrix layout change. The weight dimension of the LLM’s FFN layer, initially configured as $[8192\times 33936]$ during training on FSDP, changes to $[8192\times 8484]$ after migration to Megatron with a tensor parallelism degree of 4.
@@
After migration to Megatron, this operator exhibits significantly lower FLOPS due to smaller batch size and the unfavorable 8484 layout for Tensor Cores, which require alignment to 128 bytes. In contrast, the dimension 33936 and larger batch size on FSDP meet this alignment requirement. Following XPUTimer’s diagnosis, our infrastructure team customizes a kernel that pads 8484 to 8512. Figure 13 further illustrates the FLOPS of the same operator before migration, after migration, and post-optimization guided by XPUTimer. As shown, the operator experiences a $65.3\%$ decline in FLOPS after migration, and XPUTimer successfully facilitates the performance diagnosis. From the perspective of the training job, the overall MFU increases from $27\%$ to $36\%$, reflecting a $33.3\%$ improvement.
@@
7.4 Case-3: New Algorithms and Data
Algorithm teams continually strive to enhance model performance by modifying the LLM architecture and incorporating new training data. However, this process often introduces anomalies.
Firstly, algorithm teams generally modify position embeddings (PE), activation functions (ACT), and normalization operators (NORM), while preserving the core structure of the transformer. Table 4 illustrates the detected changes in $V_{minority}$ caused by these operator modifications during daily deployments. In this table, the parallel backbone is Megatron. The “Healthy” column represents a fully optimized training job, whereas the “-PE” column reflects changes in modifying the position embeddings. Similarly, the other columns correspond to modifications of the respective operators.
Since these modified operators are less critical and not instrumented by XPUTimer, $V_{minority}$ increases proportionally with their computational complexity. Our infrastructure team leverages XPUTimer’s detection of high $V_{minority}$ to develop targeted kernel implementations. Once optimized through techniques like kernel fusion, the job’s $V_{minority}$ returns to a normal level. In this process, XPUTimer eliminates the need for manual identification, thereby accelerating anomaly diagnosis.
Table 4: Changes in detected $V_{minority}$ and corresponding TFLOPS when different minority kernels are not optimized.

### v2

@@ 6 Evaluation
Flare consists of $11.4K$ lines of code: 2,197 lines in Python, 7,447 lines in C++, and the rest in supporting code to facilitate its usage. Beyond the automatic diagnostic workflow described above, Flare also provides rich information to assist manual optimizations, e.g., visualized distributed training timeline. In this section, we present experiments to demonstrate Flare’s effectiveness from various perspectives.
6.1 Functionality Comparison
Since most related works [jiangMegaScaleScaling, dongEnhancingLargescale, wuGREYHOUNDHunting] are closed-source or only partially open-sourced, we begin with a functionality comparison between Flare and them, followed by in-depth evaluations of Flare. Table 2 summarizes the comparison.
Firstly, in terms of tracing mechanisms, only MegaScale and Flare provide full-stack tracing. Other works, such as C4D, Greyhound, and Holmes, trace only communication or computation kernels. However, due to different design assumptions, Flare ensures backend extensibility, whereas MegaScale requires intrusive modifications to backend codebases, e.g., patching FSDP before tracing.
Secondly, in diagnostics, only Flare offers automated detection of performance regression anomalies, mainly from algorithm and infrastructure teams, enabled by the proposed micro metrics. While MegaScale supports full-stack tracing, it only provides distributed visualization for manual regression investigation. C4D, Greyhound, and Holmes diagnose only errors and fail-slows. Although Flare also detects these anomalies, it does not enhance their diagnostic mechanisms beyond providing intra-kernel inspection for fast communication hang-error diagnosis.
In a nutshell, Flare is the first diagnostic framework to combine full-stack tracing with backend extensibility and provide automated regression anomaly diagnostics.
Table 2: Functionality comparison between Flare and other existing works. “Comm.” denotes communication.
Category Features MegaScale C4D Greyhound Flare
User experience Full-stack tracing
LightCayenneBackend-extensible
Easy-to-play interfaces
Automated diagnostics
with aggregated metrics
Distributed visualization
Hang error Non-comm. hang
Comm. hang $\geq 30min$ $\geq 30min$ $\leq 5min$
Slowdown Critical kernels
Overlapping of
Comp. and Comm.
Comm. kernels
Kernel-issue stall Only GC
Less critical operations
@@ 6.2 Runtime Overhead
Figure 8: Runtime overhead in terms of latency with various models, backends, and number of GPUs.
@@
We evaluate Flare on four parallel backends: Megatron [shoeybiMegatronLMTraining], FSDP [zhaoPyTorchFSDP], DeepSpeed [rajbhandariZeROMemory], and TorchRec [ivchenkoTorchRecPyTorch]. Among these, Megatron, FSDP, and DeepSpeed are widely used for LLM training, while TorchRec is employed for training large recommendation models within Ant Group. Four models are benchmarked, spanning language, vision, and recommendation tasks: two large language models (Llama 18B and 70B), one large vision model (Llama Vision 40B), and one recommendation model (DLRM 72M).
@@
The latency overhead experiment is conducted on 1,024 H800 GPUs deployed across 128 servers with RoCE connectivity. We directly compare Flare with the original execution, as it serves as the lower bound. As shown, Flare incurs a latency overhead of $0.43\%$ for three LLM training backends and $1.02\%$ for TorchRec. Across other GPUs like A100, Flare shows consistent results.
We also compared the runtime overhead of Flare with existing works [jiangMegaScaleScaling, wuGREYHOUNDHunting]. Compared with MegaScale, Flare incurs similar runtime overhead, as both selectively trace key code segments. The main difference is that Flare avoids intruding into backend codebases, achieving better backend extensibility. Originally, Greyhound only traces start timestamps of communication kernels. We extend Greyhound’s tracing mechanism to support full-stack tracing. In this case, Greyhound incurs unacceptable overhead, reaching $35\%$ latency with Llama-8B trained on just 8 GPUs. This is because Greyhound’s tracing mechanism is tailored for fail-slows and does not cover regressions.
@@
The memory overhead experiment is conducted on two setups, which are 16 A100 GPUs on 2 nodes and 1536 H800 GPUs on 192 nodes. The testbed slightly differs from that used in the latency overhead evaluation, as we conducted the experiments on available idle GPUs due to heavy load on our training cluster. We compare Flare with Torch Full, Torch w/o Stack, Torch w/o Layout&Stack, and Flare. Torch w/o Stack refers to using the PyTorch builtin profiler with stack tracing disabled, while Torch w/o Layout&Stack further disables matrix layout tracing.
@@
Figure 9 shows the memory overhead results on 16 A100 GPUs. Flare consumes only $0.39\%$, $1.76\%$, and $2.48\%$ of memory overhead for the respective configurations of PyTorch profiler. Flare generates a maximum of $0.78\text{MB}$ of tracing logs per GPU. Besides, in a real-world Llama-20B training job on 1536 H800 GPUs, Flare generated only a $1.5\text{MB}$ tracing log per GPU. TorchRec is omitted from this experiment, as monitoring a recommendation model generates minimal logs. From the above results, Flare consistently maintains an extremely low runtime overhead in terms of both latency and memory. The lightweight selective tracing facilitates Flare’s deployment within our training cluster, serving as a diagnostic framework for diverse training jobs.
@@
Figure 9: Memory consumption of dumped logs per GPU per step using the PyTorch profiler and Flare while training a Llama-70B model on 16 A100 GPUs.
6.3 Effectiveness of Intra-kernel Inspecting
@@
We evaluate the intra-kernel inspecting mechanism on 16 A100 GPUs across two servers with RoCE connectivity. Given that most communication kernels are ring-based, this experiment focuses on evaluating ring-allreduce. We customize the training script composed solely of communication kernels, with one GPU intentionally suspended to simulate a hang error caused by communication issues.
@@
Figure 10 illustrates the pinpointing latencies for intra- and inter-server communication. The figure presents the latency results for three communication protocols [NVIDIA_NCCL_User_Guide] and cross-node configurations. As shown, Flare requires $29.4\text{\textasciitilde}309.2s$ to detect erroneous GPUs across different scenarios. Among the protocols, Flare performs best when the SIMPLE protocol is used for communication. This is because, with the SIMPLE protocol, Flare only needs to scan the first thread of each thread block to check the steps, whereas the other two protocols require scanning the entire thread block.
@@
When comparing intra-server and inter-server results, Flare performs better when the ring-allreduce operation spans multiple servers. This is because intra-server GPUs are connected via NVLink, whereas inter-server GPUs communicate through NICs. Communication kernels launched over NICs involve fewer thread blocks, as NICs have fewer internal links compared to NVLink. As a result, Flare scans fewer thread blocks for error diagnosis in inter-server scenarios.
@@
In summary, the intra-kernel inspecting mechanism can detect erroneous GPUs in a maximum of $309.2s$. Notably, as the complexity of intra-kernel inspecting is $O(1)$, these results remain constant.
@@
Figure 11: Issue distribution for a Llama-20B model trained with Megatron and 256 GPUs, including the overall CDF and the CDFs for each type of communication kernel, respectively.
@@ 6.4 Effectiveness of Issue Latency Distribution
In this experiment, we evaluate the issue latency distribution using Llama-20B running on 256 H800 GPUs across 32 servers connected via RoCE. Figure 11 illustrates the issue latency distribution for all communication kernels in the Unhealthy-GC, Unhealthy-Sync, and Healthy scenarios. In the Unhealthy-GC scenario, GC is implicitly triggered by the Python runtime. In the Unhealthy-Sync scenario, an unintended GPU synchronization call is added within the transformer block, leading to repetitive GPU synchronizations during the model’s forward pass. In the Healthy scenario, GC is efficiently managed by the parallel backend, and no unnecessary synchronizations are introduced.
As shown in the figure, the issue latency distribution patterns align with our claim in §5.2.2. The issue latency CDF of a healthy LLM training job increases linearly, whereas the issue latency CDFs for Unhealthy-GC and Unhealthy-Sync exhibit a much steeper rise. This is because the issue latencies of different ranks in the healthy scenario are solely influenced by the collective communication operator, resulting in a uniform distribution. In contrast, in the cases of Unhealthy-GC and Unhealthy-Sync, while some ranks are affected, their latencies become very short due to the delayed start of the issue time. Since both GC and GPU synchronizations span the entire model forward pass, all communication kernels are affected, as illustrated in Figure 11. Furthermore, each training process triggers GC independently, and the GC operation for a single process is more time-consuming than GPU synchronization. Consequently, the issue latency distribution for Unhealthy-GC is worse than that of Unhealthy-Sync.
@@ 7.1 Cluster-wide Deployment
Flare has been deployed and running continuously in a training cluster with 6,000 GPUs for over eight months. During this period, it is responsible for monitoring, detecting, and diagnosing training jobs for various deep learning models, especially large-scale distributed LLM training.
Table 3: Typical errors detected by Flare.
@@
7.2 Errors
Table 3 presents a subset of error and fail-slow anomalies detected by Flare. As shown in the table, Flare effectively identifies OS- and hardware-related anomalies, including crashes, and hangs. While these issues are typically conspicuous and could be detected by existing methods based on noticeable training interruptions [shoeybiMegatronLMTraining, wuGREYHOUNDHunting], Flare’s novelty lies in providing richer error information through techniques like intra-kernel inspecting. Such runtime information helps ease and accelerate the attribution process of low-level issues for operations teams.
7.3 Fail-slows and Performance Regressions
Table 4 summarizes fail-slow and regression slowdowns diagnosed by Flare using its aggregated metrics. We omit evaluation of Flare in fail-slow diagnostics, as it employs similar mechanisms to prior works [dongEnhancingLargescale, wuGREYHOUNDHunting].
To evaluate regression diagnostics, we collect 113 real-world training jobs submitted within a week in a 6,000-GPU cluster. During this period, the cluster maintained an average GPU usage of 80%. The jobs included popular LLMs, multi-modal LLMs, and recommendation models. We first diagnose them with Flare, then validate results against human-labeled ground truth. It successfully diagnosed 9 true regressions based on issue latency distribution and void percentage, with only 2 false ones. This corresponds to a false positive rate of $1.9\%$ and a true positive diagnostic accuracy of $81.8\%$.
We further investigated the 2 false positives. The first was a multi-modal LLM trained with FSDP, where input data contained images of varying resolutions. This caused imbalanced computation across ranks, leading to abnormal issue latency distribution. The second was a recommendation model using CPU-based embeddings. Its void percentage was higher than that of GPU-based ones, causing misclassification.
Although these 2 cases were falsely diagnosed as regressions, they highlight opportunities to refine Flare’s micro-metrics with historical data. For example, relaxing latency distribution thresholds for imbalanced multi-modal inputs and adjusting void percentage thresholds for CPU-based embedding models could reduce such false positives. After these refinements, Flare no longer misclassifies these job types in real-world cluster deployments.
Now, we present several typical diagnosed regressions.
Table 4: Fail-slows and regressions diagnosed by Flare, with “Details” showing training job specifics and associated MFU decline. We show the diagnosed regression anomalies by Flare in bold.
@@ 7.3 Fail-slows and Performance Regressions
Backend migration 1856 GPUs, Llama-80B, $33.3\%\downarrow$
@@
7.3.1 Case-1: Towards Stall-free Kernel Issuing
@@
Kernel-issue stalls are among the most frequent causes of slowdowns encountered in a training cluster not dedicated exclusively to a single pre-training task. While Python runtime GC is well-known and now carefully managed by the parallel backend in most cases, most encountered kernel-issue stalls arise from code introduced by algorithm teams to enhance LLM performance in downstream tasks.
@@
A typical case encountered by Flare is a training job of Llama-20B running on 256 H800 GPUs. The developer from the algorithm team mistakenly enables the timer provided by Megatron for performance profiling of several key code segments. This profiling incurs kernel-issue stalls because it requires GPU synchronizations to obtain accurate timestamps.
Although no significant regression was observed in training throughput, Flare successfully detects the abnormal issue latency distribution. After removing these unnecessary synchronizations by disabling the timer, the MFU of the training job persistently improves from $41.4\%$ to $42.5\%$, representing a $2.66\%$ increase. Such regression is minor, which could not be directly observed through macro metrics like training throughput in a large training cluster with a lot of noise. However, there is an obvious drift in micro metrics like issue latency distribution, enabling Flare to detect such obsecured regression anomalies.
@@
In addition to the above two cases, Flare also detects other kernel-issue stalls, such as unnecessary package version checking, frequent CUDA memory management within the PyTorch runtime, and others.
7.3.2 Case-2: Migration between Backends
@@
Different parallel backends are suited to varying hardware conditions. In this context, algorithm teams may migrate an LLM between backends to meet their specific demands. However, this migration process can potentially introduce regressions. A typical scenario encountered by Flare is an anomalous MFU decline when migrating a Llama-like 80B model from FSDP (1888 H800 GPUs) to Megatron (1586 H800 GPUs with a data-parallel degree of 58, pipeline-parallel degree of 8, and tensor-parallel degree of 4). This regression specifically stems from a matrix layout change. The weight dimension of the LLM’s FFN (feed forward network) layer, initially configured as $8192\times 33936]$ during training on FSDP, changes to $8192\times 8484]$ after migration to Megatron with a tensor parallelism degree of 4.
@@
After migration to Megatron, this operator exhibits significantly lower FLOPS due to smaller batch size and the unfavorable 8484 layout for Tensor Cores, which require alignment to 128 bytes. In contrast, the dimension 33936 and larger batch size on FSDP meet this alignment requirement.
While the algorithm team does not report this regression, Flare detected and routed it to our infrastructure team. Following Flare’s diagnosis, our infrastructure team customizes a kernel that pads 8484 to 8512. Figure 12 illustrates the FLOPS of the same operator before migration, after migration, and post-optimization guided by Flare. As shown, the operator experiences a $65.3\%$ decline in FLOPS after migration, and Flare successfully facilitates the performance diagnosis. From the perspective of the training job, the overall MFU increases from $27\%$ to $36\%$, reflecting a $33.3\%$ improvement.
@@
7.3.3 Case-3: New Algorithms and Data
Algorithm teams continually strive to enhance model performance by modifying the LLM architecture and incorporating new training data. However, this often introduces regressions. Firstly, algorithm teams generally modify position embeddings (PE), activation functions (ACT), and normalization operators (NORM), while preserving the core structure of the transformer. Table 5 illustrates the detected changes in $V_{minority}$ caused by these operator modifications during daily deployments. In this table, the parallel backend is Megatron. The Healthy column represents a fully optimized training job, whereas the -PE column reflects changes in modifying the position embeddings. Similarly, the other columns correspond to modifications of the respective operators.
Since these modified operators are less critical and not instrumented by Flare, $V_{minority}$ increases proportionally with their computational complexity. Our infrastructure team leverages Flare’s detection of high $V_{minority}$ to develop targeted kernel implementations. Once optimized through techniques like kernel fusion, the job’s $V_{minority}$ returns to a normal level. In this process, Flare eliminates the need for manual identification, thereby accelerating anomaly diagnosis.
Secondly, newly filtered data is continuously incorporated into the LLM for training. Flare has successfully diagnosed anomalies arising from variance in the training data. In one specific case, the algorithm team attempts to train Llama-80B with data containing a sequence length of 64k, while the original training script is for sequence lengths of 4k. Flare identifies a significant anomalous decline in MFU ($41\%$) on 512 H800 GPUs, accompanied by an increase in $V_{inter}$.
After routing this anomaly with the root cause identified in the dataloader to the algorithm team, they validated that the regression was due to the attention mask generation process within the dataloader. When the sequence length is short, the latency incurred by mask generation is minimal. However, the complexity of mask generation scales as $O(L^{2})$, where $L$ represents the sequence length. As a result, the dataloader experiences extremely poor performance when the sequence length increases to 64k.
Table 5: Changes in detected $V_{minority}$ and normalized TFLOPS when different minority kernels are not optimized.